Integrals Giving Inverse Trigonometric Functions


∫ dy/√(25+9y²)


1
Expert's answer
2021-04-15T16:53:59-0400

Solution.

"\int\frac{dy}{\sqrt{25+9y^2}}"

Make a substitution: "u=\frac{3y}{5}," then "\frac{du}{dy}=\frac{3}{5}," since "dy=\frac{5}{3}du."

So, "\int\frac{dy}{\sqrt{25+9y^2}}=\frac{1}{3}\int \frac{du}{\sqrt{u^2+1}}."

Make a substitution: "u=\tan{v}," then "v=\arctan{u}," since "du=\sec^2{v}dv."

So,

"\frac{1}{3}\int \frac{du}{\sqrt{u^2+1}}=\int \frac{\sec^2{v}dv}{\sqrt{\tan^2{v}+1}}=
\int \sec{v}dv=\newline
=\int \frac{\sec{v}(\tan{v}+\sec{v})}{\tan{v}+\sec{v}}dv."

Make a substitution: "w=\tan{v}+\sec{v}," then "\frac{dw}{dv}=\sec{v}\tan{v}+\sec^2{v}," since "dv=\frac{1}{\sec{v}\tan{v}+\sec^2{v}}dw."

So,

"\int \frac{\sec{v}(\tan{v}+\sec{v})}{\tan{v}+\sec{v}}dv=\int \frac{1}{w}dw=\ln|w|+C=
\newline =\ln |\tan{v}+\sec{v}|+C=\ln |\sqrt{u^2+1}+u|+C=\frac{1}{3}\ln |\sqrt{9y^2+25}+3y|+C,"

where "C" is some constant.

Answer. "\int\frac{dy}{\sqrt{25+9y^2}}=\frac{1}{3}\ln |\sqrt{9y^2+25}+3y|+C."